OpenStudy (anonymous):

Solve 12 x 3 - 3(2 x 3 + 4x -1) - 5x + 7

OpenStudy (anonymous):

@tcarroll010 can u solve it?

OpenStudy (anonymous):

@dumbcow ?

OpenStudy (dumbcow):

wheres the equal sign ?

OpenStudy (anonymous):

at the end of seven

OpenStudy (anonymous):

its 12 x 3 - 3(2 x 3 + 4x -1) - 5x + 7 = ?

OpenStudy (dumbcow):

??? wait are those "x's" variables or is this just a computation problem?

OpenStudy (anonymous):

the x's are multiplication signs sorry

OpenStudy (anonymous):

hold on let me write it correctly

OpenStudy (dumbcow):

in that case, use Order of Operations or put it in calculator :)

OpenStudy (anonymous):

12(3) - 3(2(3) + 4x -1) - 5x + 7 that the problem

OpenStudy (dumbcow):

36 - 3(6 +4x-1) -5x +7 36 - 18 -12x +3 -5x +7 28 -17x

OpenStudy (anonymous):

then u divide 28 and -17?

OpenStudy (dumbcow):

i dunno ....does it equal 0

OpenStudy (anonymous):

noo it equals -1.64705

OpenStudy (dumbcow):

28 -17x = 0 28 = 17x 28/17 = x 1.64705 = x